OpenStudy (anonymous):

Use Property 1 for radicals to simplify the following radical expression as much as possible. Assume the variable represents a positive number.

OpenStudy (anonymous):

\[2a \sqrt[3]{8a^4}\]

hartnn (hartnn):

factor 8

OpenStudy (anonymous):

4,2

hartnn (hartnn):

factor it further

OpenStudy (anonymous):

2,2

hartnn (hartnn):

so 8 = 2*2*2 = 2^3 right ? \(x^4 = x \times x \times x \times x = x^3 \times x\) got this ?

OpenStudy (anonymous):

I think so...

OpenStudy (anonymous):

Oh, yea i got it

hartnn (hartnn):

i am paring 3 constants and variables this time (instead of 2 which i did last time) because in this case, its cube root, instead of square root

OpenStudy (anonymous):

ok, I'm ready

hartnn (hartnn):

\(\large \sqrt[3]{\heartsuit ^3}= \heartsuit\)

hartnn (hartnn):

\(\large \sqrt[3]{\heartsuit ^3}= (\heartsuit)^{3/3}= \heartsuit \)

OpenStudy (anonymous):

2

hartnn (hartnn):

\(\Large 2a \sqrt[3]{8a^4} = 2a \sqrt[3]{8} \sqrt[3]{a}\sqrt [3]{a^3}=..\)

hartnn (hartnn):

\(\Large 2a \sqrt[3]{8a^4} = 2a \sqrt[3]{2^3} \sqrt[3]{a}\sqrt [3]{a^3}=..\)

OpenStudy (anonymous):

I know the 3s cancel out

hartnn (hartnn):

so \(\Large 2a \sqrt[3]{8a^4} = 2a \times 2 \times \sqrt[3]{a}\times a=..\) got this ?

OpenStudy (anonymous):

yup

hartnn (hartnn):

simplify

OpenStudy (anonymous):

4a?

hartnn (hartnn):

2a times 2a = 4a^2 \(\Large 4a^2 \times \sqrt[3]a\) thats it!

OpenStudy (anonymous):

Ohhhhhh! Thanks

hartnn (hartnn):

welcome ^_^